OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 2i

OpenStudy (anonymous):

1 - 4i -3 5 1 + 4i

jimthompson5910 (jim_thompson5910):

hint: (a+bi)*(a-bi) = a^2 + b^2

OpenStudy (anonymous):

Last one ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what are a and b in this case

jimthompson5910 (jim_thompson5910):

a+bi 1+2i a = ? b = ?

OpenStudy (anonymous):

1-2i ?

OpenStudy (anonymous):

a = 1 b = 2i

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

a = 1 b = 2

jimthompson5910 (jim_thompson5910):

(a+bi)*(a-bi) = a^2 + b^2 (1+2i)*(1-2i) = 1^2 + 2^2 ... plug in a = 1 and b = 2 (1+2i)*(1-2i) = ??

OpenStudy (anonymous):

1 - 2i +2i - 4i^2

jimthompson5910 (jim_thompson5910):

the -2i and 2i terms add to 0i or just 0

OpenStudy (anonymous):

so 1-4i^2

jimthompson5910 (jim_thompson5910):

you now have 1 - 4i^2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

i^2 = -1 1 - 4i^2 turns into 1 - 4(-1)

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

correct